A rectangle has length 2 cm longer than its width. If the rectangle has perimeter 52 cm, find its width.

Answer by Boreal(15235)

width=x
length=x+2
half perimeter is L*W=26 (half of 52)
x+x+2=26
2x+2=26
x=12 cm width, the answer
x+2=14 cm length